OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}[\frac{ e ^{tanx}-e^x }{ tanx - x }]\]

OpenStudy (anonymous):

so just use l'hopitals

OpenStudy (anonymous):

try that and then i can fix if you have more problems

OpenStudy (anonymous):

i tried L-Hospital..it is werid...

OpenStudy (anonymous):

d(e^tanx -e^x)/dx e^tanx (sec^2 x) -e^x

OpenStudy (anonymous):

d(tanx -x)/dx sec^2 x -1

OpenStudy (anonymous):

^ yup.

Parth (parthkohli):

L'Hospital's, Chain Rule and Quotient Rule! I'd die doing this question O_O

OpenStudy (anonymous):

But lim x--->0 sec^2 x-1 =0

OpenStudy (anonymous):

Again use L'Hospital's Rule ???

OpenStudy (anonymous):

sec^2 x-1 = tan^2 x

OpenStudy (experimentx):

try using expressing in this form ... sorry don't have much time http://www.wolframalpha.com/input/?i=lim+x-%3E0+%28e^x-1%29%2Fx

OpenStudy (experimentx):

this is interesting